if the 5th term & 8th term =30 & if the 25th term is 3 times the 9th term . Find the ap
Answer by richwmiller(17219)

a + 4d+ a +7d=30,
2a+11d=30,
a + 24d=3*(a + 8d)
a + 24d=3a + 24d
a = 0, d = 30/11
an=(n-1)*(30/11)
a5=(4)*(30/11)=120/11
a8=(7)*(30/11)=210/11
a9=8*30/11=240/11
a25=24*30/11=720/11